1)A 1.00 liter solution contains 0.55 moles hydrofluoric acid and 0.42 moles sodium fluoride .

If 0.21 moles of hydroiodic acid are added to this system, indicate whether the following statements are true or false.

(Assume that the volume does not change upon the addition of hydroiodic acid.)

_______TrueFalseA. The number of moles of HF will increase.

_______TrueFalseB. The number of moles of F- will remain the same.

_______TrueFalseC. The equilibrium concentration of H3O+ will decrease.

_______TrueFalseD. The pH will decrease.

_______TrueFalseE. The ratio of [HF] / [F-] will remain the same.

2)A 1.00 liter solution contains 0.53 M ammonia and 0.41 M ammonium bromide.

If 0.2 moles of hydroiodic acid are added to this system,
indicate whether the following statements are true or false.

(Assume that the volume does not change upon the addition of hydroiodic acid.)

_______TrueFalse A. The number of moles of NH3 will decrease.

_______TrueFalse B. The number of moles of NH4+ will increase.

_______TrueFalse C. The equilibrium concentration of H3O+ will decrease.

_______TrueFalse D. The pH will increase.

_______TrueFalse E. The ratio of [NH3] / [NH4+] will remain the same.
Answer #1

1)

We have Henderson's Equation, pH = pKa + log [ Salt ] / [acid ]

pH = pKa + log [ NaF ] / [ HF ]

pH = 3.17 + log ( 0.42 / 0.55 )

pH = 3.17 - 0.117

pH = 3.05

We have relation, pH = - log [ H3O + ]

[ H3O + ] = 10 -pH

[ H3O + ] = 10 - 3.05 = 8.85 10 -04 M

Ratio [ HF ] / [ F - ] = 0.55 / 0.42 = 1.31

HI is a strong halo acid. Consider its reaction with buffer solution.

F - + HI HF + I -

Let's use ICE table.

moles F - + HI HF
I 0.42 0.21 0.55
C - 0.21 -0.21 +0.21
E 0.21 0.00 0.76

We have Henderson's Equation, pH = pKa + log [ Salt ] / [acid ]

pH = pKa + log [ NaF ] / [ HF ]

pH = 3.17 + log ( 0.21 / 0.76 )

pH = 3.17 - 0.558

pH = 2.61

We have relation, pH = - log [ H3O + ]

[ H3O + ] = 10 -pH

[ H3O + ] = 10 - 2.61 = 2.44 10 -03 M

Ratio [ HF ] / [ F - ] = 0.76 / 0.21 = 3.62

From above calculations we can answer given statements.

True  A. The number of moles of HF will increase.

False B. The number of moles of F- will remain the same.

False C. The equilibrium concentration of H3O+ will decrease.

True D. The pH will decrease.

False E. The ratio of [HF] / [F-] will remain the same.

QUESTION 2

We have Henderson's Equation, pH = pKa + log [ Salt ] / [acid ]

pH = pKa + log [NH 3 ] / [ NH 4+ ]

pH = 9.245 + log ( 0.53/ 0.41 )

pH =9.245 + 0.111

pH = 9.356

We have relation, pH = - log [ H3O + ]

[ H3O + ] = 10 -pH

[ H3O + ] = 10 - 9.356 =4.40 10 -10 M

Ratio [NH 3 ] / [ NH 4+ ] = 0.53 / 0.41 = 1.29

Consider reaction of HI with buffer solution.

NH 3 + HI NH 4+ + I -

Let's use ICE table.

moles NH 3 + HI NH 4+
I 0.53 0.2 0.41
C -0.2 -0.2 +0.2
E 0.51 0.0 0.43

We have Henderson's Equation, pH = pKa + log [ Salt ] / [acid ]

pH = pKa + log [NH 3 ] / [ NH 4+ ]

pH = 9.245 + log ( 0.51/ 0.43 )

pH =9.245 + 0.074

pH = 9.319

We have relation, pH = - log [ H3O + ]

[ H3O + ] = 10 -pH

[ H3O + ] = 10 - 9.319 =4.80 10 -10 M

Ratio [NH 3 ] / [ NH 4+ ] = 0.51 / 0.43 = 1.19

From above calculations we can answer given statements.

True A. The number of moles of NH3 will decrease.

True B. The number of moles of NH4+ will increase.

False C. The equilibrium concentration of H3O+ will decrease.

False D. The pH will increase.

False E. The ratio of [NH3] / [NH4+] will remain the same.